OpenStudy (anonymous):

cos3x over cosx=1-4sin^2 x verify the identities

OpenStudy (ujjwal):

\(\cos3x=4\cos^3x-3\cos x\)

OpenStudy (campbell_st):

well you need to trig expansion for cos(3x) make it cos(2x + x) so you are looking at \[\frac{\cos(2x)\cos(x) - \sin(2x)(sinx)}{\cos(x)}\] and the simplfy from here... substitute sin(2x) = 2sin(x)cos(x) and \[\cos(2x) = \cos^2(x) - \sin^2(x)\] then its just the manipulation into your desired answer

OpenStudy (anonymous):

Sorry but I still don't get it I'm completely lost

OpenStudy (campbell_st):

ok.... do you know about cos(2x) and its expansion or cos(a + b) and its expansion...

OpenStudy (anonymous):

Nope

OpenStudy (campbell_st):

wow... then it is virtually impossible for you to do the question using your trig skills so use the first piece of information you have \[\cos(3x) = 4\cos^3(x) - 3\cos(x)\] substitute it into your equation and you get \[\frac{4\cos^3(x) - 3\cos(x)}{\cos(x)}\] which can be written as \[\frac{4\cos^3(x)}{\cos(x)} - \frac{3\cos(x)}{\cos(x)}\] do you think you could simplify the above fractions by removing the common factor?